If P = 7y + 2xy - 6x and Q = 5 - 7y, find the value of 3P - 4Q. Please help I am very confused on this one. Please Help Me

Answer by checkley71(8403)

P=7Y+2XY-6X
Q=5-7Y THEN:
3P=3(7Y+2XY-6X)=21Y+6XY-18X
4Q=4(5-7Y)=20-28Y
NOW SUBTRACT 4Q(20-28Y) FROM 3P(21Y+6XY-18X)
21Y+6XY-18X-(20-28Y)
21Y+6XY-18X-20+28Y
49Y+6XY-18X-20 ANSWER.